# Are Large Language Models Consistent over Value-laden Questions?

### v1

@@ Abstract
Large language models (LLMs) appear to bias their survey answers toward certain values. Nonetheless, some argue that LLMs are too inconsistent to simulate particular values. Are they? To answer, we first define value consistency as the similarity of answers across (1) paraphrases of one question, (2) related questions under one topic, (3) multiple-choice and open-ended use-cases of one question, and (4) multilingual translations of a question to English, Chinese, German, and Japanese. We apply these measures to a few large ($>=34b$), open LLMs including llama-3, as well as gpt-4o, using eight thousand questions spanning more than 300 topics. Unlike prior work, we find that models are relatively consistent across paraphrases, use-cases, translations, and within a topic. Still, some inconsistencies remain. Models are more consistent on uncontroversial topics (e.g., in the U.S., Thanksgiving) than on controversial ones (euthanasia). Base models are both more consistent compared to fine-tuned models and are uniform in their consistency across topics, while fine-tuned models are more inconsistent about some topics (euthanasia) than others (women’s rights) like our human subjects (n=165).
@@ 1 Introduction
Figure 1: Similar to our human participants (n=84), chat model are inconsistent (change their answers) on topics like euthanasia and religious freedom but they are consistent on topics like women’s rights and income inequality. This is less the case for base models like llama3-base. To measure such topic inconsistency, we prompted models with similar questions about a specific topic, measuring the distance between answers using a variant of the Jensen-Shannon divergence, the D-dimensional divergence (§3.2). Shown here are the two topics with the highest and lowest topic inconsistency across models in English on U.S.-based topics; other languages and topics reported elsewhere.
@@
We detail an unsupervised method to gauge the consistency of models’ expressed behavior as a means to quantify what values models have. To do so, we formalize a number of desirable measures of value consistency, assuming that the values latent in an answer to a particular question remain reasonably consistent across (1) paraphrases, (2) multiple-choice and open-ended use-cases, (3) multilingual translations, and (4) across similar questions within a given topic3). While these measures may be used for consistency more broadly, we call them measures of value consistency here as they operate in explicitly value-laden domains. In order to apply these measures, we introduce a novel dataset, ValueConsistency, containing more than 8k questions over 300 topics and four languages (§4).
Unlike prior work, we investigate both controversial and uncontroversial topics, compare base models and fine-tuned models, generate country-specific topics, and study models’ consistency over translations. Via extensive analyses, we find the following: (1) Contrary to our expectations, large models are reasonably consistent over our measures, being as or more consistent than our human participants (n=165) (Fig. 4). (2) Across measures, models are more consistent over less controversial questions (Fig. 5). (3) Base models are more consistent compared to their fine-tuned counterparts (Fig. 3). (4) Fine-tuned models, like our human participants, are more consistent on some topics than others; base models are equally consistent (Fig. 6).
@@ 2.3 Model Steerability
A variety of scholars have attempted to steer models to particular values, especially to align the distribution of a model’s responses over a domain to the distribution of some group (e.g. “Answer like a Democrat”) (Santurkar et al., 2023) or persona (Shu et al., 2024; Liu et al., 2024), although a few note that prior survey responses, more than any particular group label, are better predictors of future responses (Zhao et al., 2023; Hwang et al., 2023; Li et al., 2023a). Wang et al. (2024a) are critical of this space, finding that LLMs tend toward erroneous portrayal of identity groups.
@@ 3.2 Distance between Answers
Now, eq. 1 compares just two distributions. Given a list of distributions we thus calculate the Jensen-Shannon centroid, the distribution closest to all given distributions (Nielsen, 2020).
@@
We define the d-dimensional Jensen-Shannon divergence (D-D div., for short) which is the average divergence between each distribution and their centroid (eq. 2):
@@ 3.3 Consistency Measures
We lay out a framework for assessing values, defining a number of existing and new measures of consistency. We formalize them in Tab. 1 and further explain each in §A.3.
@@ 4 Constructing ValueConsistency
Instead of relying on existing datasets of controversial topics such as surveys (Santurkar et al., 2023), we sought to provide an extensible, and largely unsupervised, method to generate value-relevant questions. Indeed, prior work has used LLMs to systematically generate, with reliable filtering, the content of datasets for social NLP Ziems et al. (2023a); Scherrer et al. (2023); Fränken et al. (2023); Gandhi et al. (2023). We thus introduce ValueConsistency, a dataset of more than 8000 questions across more than 300 topics. Tab. 2 breaks down our questions by category and Tab. 6 lists a few example topics. [^3] Our data and code will be available under the CC-BY 4.0 license here after reviewing
In particular, we generated topics, questions relevant to those topics, answers to those questions with their associated stance toward a topic (e.g., “yes” to “do you like cats” indicates support for cats), and paraphrases for those questions. See Fig. 2. We prompted for controversial topics in the United States in English, translating them to Chinese, German, and Japanese using gpt-4-0613. We did the same for topics in each subsequent country and language, but for the rest only translated to English. [^4] We recognize that countries are not cultural monoliths and culture is not simply nationality (Adilazuarda et al., 2024). Our questions do not capture all cultural nuances. We chose these languages because they are common, geographically diverse, and we could find a large, pre-trained alignment-tuned model performant on them. In addition to controversial topics, we also compared against generated uncontroversial topics as a baseline.
@@ 5 Experiment Setup
Tab. 3 shows the large models we queried and in which of Chinese, Japanese, English, German. We followed standard prompting best practices. For the multiple-choice use-case we gathered models’ option-token log probabilities (Wang et al., 2024c) (e.g. “A”, “B”, etc.). For the open-ended use-case, we used llama3 to detect the stance and classify each model response. Further details in §C.
Table 3: Models. We refer to models by their abbreviated “fine-tuned” and “base” names. cmd-r is Command R from Cohere. More info in §C.
@@
llama2 llama2-base 70b eng, chi,
ger, jpn
llama3 llama3-base 70b eng, chi,
ger, jpn
cmd-R 35b eng, chi,
ger, jpn
@@
Human Annotation
We administered our survey to human participants, but only on controversial U.S.-based topics in English. Our institution’s IRB approved this study. We paid participants more than the federal minimum. For topic consistency (n=84), we asked each unique participant multiple related questions about one topic. For paraphrase consistency (n=81), we asked each unique participant one unique question per topic and all paraphrases of that question. We compute participants’ consistency using the D-D divergence, and average consistency between them. More info in §C.
@@ 6 Results
Figure 3: Base models are more consistently consistent unlike chat models and human participants. On the x-axis is each topic ordered by least to most consistent in English on U.S.-based topics. Each colored bar shows either the topic consistency (top plots) or paraphrase consistency (bottom plots). Both fine-tuned models and human participants (n=84 for topic, n=81 for paraphrase) show a greater spread than base models. Error bars show 95% bootstrapped confidence intervals.The dashed line shows the upper limit of .46 for our measure of inconsistency, the D-D divergence (§3.2, §A.2).
Figure 4: Models are relatively consistent across our measures. They are as or more consistent than our human participants (n=81 for paraphrase and n=84 for topic consistency, §5). In these plots we only compare topics for the U.S. in English (except in multilingual consistency, where we compare across up to all of {eng, chi, ger, jpn}). Error bars show 95% bootstrapped confidence intervals.
@@ 6.1 Consistency across topics
Within each model, we compared measures of consistency across topics. Fine-tuned models are much more inconsistent than base models when compared by topic. For example, llama3-base is about 60% more topic consistent than llama3. See Fig. 3. Namely, llama3 significantly more inconsistent on euthanasia with a mean score of about .4 than it is on women’s rights with a mean of score of 0 while llama3-base is roughly as consistent in both cases (ascoring bout .2 and .1, respectively). See Fig. 1. In both topic and paraphrase consistency, fine-tuned models are more similar to our human participants in being inconsistently inconsistent (Fig. 3). For example, the mean topic inconsistency for our human respondents was .29 with a max of .44 and a min of 0, akin to the mean topic consistency of llama3 of .19 with a max of .45 and min of 0 compared to the mean for llama3-base of .12 with a max of .20 and min of .07.
@@ 6.2 Consistency by {un}controversial
Figure 7: Chat models are much less consistent on topics like euthanasia than they are for topics like women’s rights while base models are similarly consistent. Shown are the four topics with the highest (top row) and lowest (bottom row) topic inconsistency across models and human participants (n=84) in English on U.S.-based topics. Questions for each topic shown in Tab. 9 and 10.
@@
Comparing alignment fine-tuned models with their base model equivalents (Tab. 3), Fig. 6 shows that base models are more consistent, especially on topic consistency. For example, llama3 is about 60% more topic consistent than llama3-base. While llama3 is about 33% less paraphrase consistent than llama3-base, all other chat models are more paraphrase consistent than their base models.
@@ 6.4 Consistency by use-case
We find that models are generally somewhat less consistent in the open-ended use-case than in the multiple-choice use-case (§3). This is more pronounced for yi and stability which are 27% and 57% more topic consistent on multiple-choice as shown in Fig. 8. Only llama2 is less topic consistent on multiple-choice with a reduction of 20%. Note that we use llama3 to judge the stance of the open-ended generations, and we find that it achieves substantial agreement with claude-3-opus and gpt-4o, with a median Fleiss’s Kappa of 0.7. (See Fig. 11.)
Figure 8: Chat models are somewhat less consistent in the open-ended use-case than in the multiple-choice use-case. We prompt gpt-4o, llama2, llama3 with U.S. topics and cmd-r, yi, and stability with German, Chinese, and Japanese topics, each in their respective dominant languages. We use llama3 to judge the stance of the open-ended generations. Error bars show 95% bootstrapped confidence intervals.
@@ 6.5 Can models be steered to certain values?
Scholars often care about not just which values models express but also to which they are sensitive. Here we study whether models can be steered to answer in line with Schwartz’s values (Schwartz, 1992) as a proxy for value steerability more generally. We choose Schwartz’s values because previous work has shown mixed results as to whether LLMs are steerable to them (Zhang et al., 2023; Yao et al., 2023; Fischer et al., 2023).
To determine whether prompting with certain value-words has any effect on models, we must first determine whether models can disambiguate between different values when prompted. To do so, we prompted models with the questionnaire used to cluster and create Schwartz’s 11 values, the Portrait Values Questionnaire (PVQ-21). We then tested whether appending the name of each value (e.g. “universalism”) had a larger effect on the model response as compared to values unrelated to the question. (§A.4 offers a formal treatment. See §D.2 for an example.)
We ask: which value was the most influential, the relevant value or an unrelated value? A rank of 0 indicates all of the unrelated values had a bigger effect than the related value while a rank of 11 (for the 12 values) means that the relevant value had a bigger effect than the unrelated values. While we would expect high rankings—high “steerability”—instead we find that unrelated values are more influential than relevant ones (Fig. 9). This means that the models were not steerable to these values. We found similar results across the languages we tested, although the PVQ-21 was not available in Japanese (Schwartz, 2021).
@@ 7 Discussion
Prior work has argued that models either do (Durmus et al., 2024; Santurkar et al., 2023) or do not (Röttger et al., 2024; Shu et al., 2024) hold certain values. So: Are LLMs consistent over value-laden questions? While the answer is more yes than no, our findings show that the underlying complexity cannot be captured by a binary answer.
Indeed, unlike prior work (Röttger et al., 2024; Shu et al., 2024), we have found that large models ($>=34b$) are relatively consistent across our measures, performing on par with human participants on topic and paraphrase consistency (Fig. 4). Nonetheless, models’ consistency is not uniform.
In general, base models are more consistent than their fine-tuned counterparts (Fig. 5). Moreover, base models are more consistently consistent than fine-tuned ones. For example, llama3, like our human participants, is very consistent on women’s rights but very inconsistent on euthanasia while llama3-base does not exhibit such patterns (Fig. 3). Models are more consistent over uncontroversial questions than controversial ones (Fig. 6). In addition, we measure how well models can be steered to particular values (§6.5), showing that models cannot be steered using a common set of values (Fig. 9).

### v2

@@ Abstract
Large language models (LLMs) appear to bias their survey answers toward certain values. Nonetheless, some argue that LLMs are too inconsistent to simulate particular values. Are they? To answer, we first define value consistency as the similarity of answers across (1) paraphrases of one question, (2) related questions under one topic, (3) multiple-choice and open-ended use-cases of one question, and (4) multilingual translations of a question to English, Chinese, German, and Japanese. We apply these measures to small and large, open LLMs including llama-3, as well as gpt-4o, using 8,000 questions spanning more than 300 topics. Unlike prior work, we find that models are relatively consistent across paraphrases, use-cases, translations, and within a topic. Still, some inconsistencies remain. Models are more consistent on uncontroversial topics (e.g., in the U.S., "Thanksgiving") than on controversial ones ("euthanasia"). Base models are both more consistent compared to fine-tuned models and are uniform in their consistency across topics, while fine-tuned models are more inconsistent about some topics ("euthanasia") than others ("women’s rights") like our human subjects (n=165).
@@ 1 Introduction
Figure 1: Similar to our human participants (n=84), chat models are inconsistent (change their answers) on topics like euthanasia and religious freedom but they are consistent on topics like women’s rights and income inequality. This is less the case for base models like llama3-base. To measure such topic inconsistency, we prompted models with similar questions about a specific topic, measuring the distance between answers using a variant of the Jensen-Shannon divergence, the D-dimensional divergence (§3.2). Shown here are the two topics with the highest and lowest topic inconsistency across models in English on U.S.-based topics; other languages and topics reported elsewhere.
@@
We detail an unsupervised method to gauge the consistency of models’ expressed behavior as a means to quantify what values models have. To do so, we formalize a number of desirable measures of value consistency, assuming that the values latent in an answer to a particular question remain reasonably consistent across (1) paraphrases, (2) multiple-choice and open-ended use-cases, (3) multilingual translations, and (4) across similar questions within a given topic3). While these measures may be used for consistency more broadly, we call them measures of value consistency here as they operate in explicitly value-laden domains. In order to apply these measures, we introduce a novel dataset, ValueConsistency, containing more than 8k questions over 300 topics and across four languages (§4).
Unlike prior work, we investigate both controversial and uncontroversial topics, compare base models and fine-tuned models, generate country-specific topics, and study models’ consistency over translations. Via extensive analyses, we find the following: (1) Contrary to our expectations, large models are reasonably consistent over our measures, being as or more consistent than our human participants (n=165) (Fig. 4). (2) Across measures, models are more consistent over less controversial questions (Fig. 5). (3) Base models are more consistent compared to their fine-tuned counterparts (Fig. LABEL:fig:models-topicwise-topic-consistency-hist-base). (4) Fine-tuned models, like our human participants, are more consistent on some topics than others; base models are equally consistent (Fig. 6).
@@ 2.3 Model Steerability
A variety of scholars have attempted to steer models to particular values, especially to align the distribution of a model’s responses over a domain to the distribution of some group (e.g. “Answer like a Democrat”) (Santurkar et al., 2023) or persona (Shu et al., 2023; Liu et al., 2024), although a few note that prior survey responses, more than any particular group label, are better predictors of future responses (Zhao et al., 2023; Hwang et al., 2023; Li et al., 2023a). Wang et al. (2024a) are critical of this space, finding that LLMs tend toward erroneous portrayal of identity groups.
@@ 3.2 Distance between Answers
Now, eq. 1 compares just two distributions. Given a list of distributions we thus calculate the Jensen-Shannon centroid, the distribution which minimizes the average JS divergence with other distributions (Nielsen, 2020).
@@
We (re)define the d-dimensional Jensen-Shannon divergence (D-D div., for short) which is the average divergence between each distribution and their centroid (eq. 2):
@@
We make no claim as to the novelty of the D-D divergence, which is very similar to the generalized JSD (eq. 6) introduced by Sibson (1969) which uses the average distribution, an approximate centroid, instead of the actual centroid, $\mathcal{C}^{*}$. Likewise, it is similar to the divergence used by Scherrer et al. (2023): just take the mean of all of the pairwise divergences (eq. 7).
@@ 3.3 Consistency Measures
We lay out a framework for assessing values, defining a number of existing and new measures of consistency (see Tab. 1).
Paraphrase Consistency
Differently expressed but semantically equivalent statements have long been a standard to judge NLP systems against (Jurafsky and Martin, 2024). Just so with values. For example, Do you think that euthanasia is morally acceptable? and In your view, is euthanasia morally acceptable? should yield the same answer (“yes” or “no” but not both).
Topic Consistency
Similar questions—those concerning the same topic—should likewise have similar answers. For example, answering “yes” to the question Do you think that euthanasia is morally acceptable? often entails the same to Do you believe that euthanasia should be legalized? Nonetheless, expect less topic consistency than paraphrase consistency; e.g., one might morally, but not legally, oppose euthanasia.
Use-case (Task) Consistency
Similar to survey design (Krosnick, 2018), prior work has used forced-choice, multiple-choice paradigms to interrogate models (Santurkar et al., 2023). These set-ups may not generalize (Röttger et al., 2024). Similarly, we compare answers to multiple-choice and open-ended questions. For example, the multiple-choice answer of “yes” (support for euthanasia) to the question, Do you think that euthanasia is morally acceptable?, ought to imply that open-ended arguments about that same question have an equivalently supporting stance.
Multilingual Consistency
A person fluent in multiple languages will answer translations of the same question similarly. Here we expect some noise due to the imperfection of translation. We compare between each of the languages in which a model can respond. As explained in §4, we generate questions pertinent to a specific country. Thus, here we keep the country constant. We also compare only the multiple-choice tasks.
@@ 4 Constructing ValueConsistency
Instead of relying on existing datasets of controversial topics such as surveys (Santurkar et al., 2023), we sought to provide an extensible, and largely unsupervised, method to generate value-relevant questions. Indeed, prior work has used LLMs to systematically generate, with reliable filtering, the content of datasets for social NLP Ziems et al. (2023a); Scherrer et al. (2023); Fränken et al. (2023); Gandhi et al. (2023). We thus introduce ValueConsistency, a dataset of more than 8000 questions across more than 300 topics. Tab. 2 breaks down our questions by category and Tab. 7 lists a few example topics. [^4] ValueConsistency is available under the MIT license here: https://huggingface.co/datasets/jlcmoore/ValueConsistency
In particular, we generated topics, questions relevant to those topics, answers to those questions with their associated stance toward a topic (e.g., “yes” to “do you like cats” indicates support for cats), and paraphrases for those questions. See Fig. 2. We prompted for controversial topics in the United States in English, translating them to Chinese, German, and Japanese using gpt-4-0613. We did the same for topics in each subsequent country and language, but for the rest only translated to English. [^5] We recognize that countries are not cultural monoliths and culture is not simply nationality (Adilazuarda et al., 2024). Our questions do not capture all cultural nuances. We chose these languages because they are common, geographically diverse, and we could find a large, pre-trained alignment-tuned model performant on them. In addition to controversial topics, we also compared against generated uncontroversial topics as a baseline. [^6] Our code is available under the MIT license here: https://github.com/jlcmoore/ValueConsistency
@@
We also recruited human annotators to validate the our dataset in non-English languages, targeting users either fluent in English and each of {Chinese, German, Japanese}. To judge the quality of our paraphrases, we presented (n=546) annotators with paraphrases of a single question in the original language, asking them in English to respond “yes” or “no” to the question, Are all these questions asking the same thing? Annotators rated paraphrases as significantly more equivalent than chance. For the controversiality of our questions, we similarly presented (n=562) annotators with a set of questions about one topic and asked them in English to respond “yes” or “no” to the question, How controversial are these questions? Annotators rated questions gpt-4 judged as controversial as significantly more controversial than questions judged as uncontroversial. (See Tab. 4.)
@@ 5 Experiment Setup
Tab. 3 shows the models we queried and in which of Chinese, Japanese, English, German. We followed standard prompting best practices. For the multiple-choice use-case we gathered models’ option-token log probabilities (Wang et al., 2024c) (e.g. “A”, “B”, etc.). Unlike the larger models (and with the exception of llama3-8b) smaller models ($<34b$) we tested, such llama2-7b, displayed an order bias. For the open-ended use-case, we used llama3 to detect the stance and classify each model response. Further details in §C.
Table 3: Models. We refer to models by their abbreviated “fine-tuned” and “base” names. cmd-r is Command R from Cohere. “All” refers to: eng, chi, ger, jpn. More info in §C.
@@
llama2 llama2-base 70b All
llama2-7b llama2-base-7b 7b All
llama3 llama3-base 70b All
llama3-8b llama3-base-8b 8b All
cmd-R 35b All
@@
Human Subjects
We administered our survey to human participants, but only on controversial U.S.-based topics in English. Our institution’s IRB approved this study. We paid participants more than the federal minimum. For topic consistency (n=84), we asked each unique participant multiple related questions about one topic. For paraphrase consistency (n=81), we asked each unique participant one unique question per topic and all paraphrases of that question. We compute participants’ consistency using the D-D divergence, and average consistency between them. We used a within-subjects design: finding how consistent a single person was across a set of questions and then averaging that across all participants. More info in §C.
@@ 6 Results
Figure 3: Base models are more consistently consistent unlike chat models and human participants. On the x-axis is each topic ordered by least to most consistent in English on U.S.-based topics. Each colored bar shows either the topic consistency (top plots) or paraphrase consistency (bottom plots). Both fine-tuned models and human participants (n=84 for topic, n=81 for paraphrase) show a greater spread than base models. Error bars show 95% bootstrapped confidence intervals.The dashed line shows the upper limit of .46 for our measure of inconsistency, the D-D divergence (§3.2, §A.2).fig:models-topicwise-topic-consistency-hist-base
Figure 4: Models are relatively consistent across our measures. They are as or more consistent than our human participants (n=81 for paraphrase and n=84 for topic consistency, §5). In these plots we only compare topics for the U.S. in English (except in multilingual consistency, where we compare across up to all of {eng, chi, ger, jpn}). Error bars show 95% bootstrapped confidence intervals.The dashed line shows the upper limit of .46 for our measure of inconsistency, the D-D divergence (§3.2, §A.2).
@@ 6.1 Consistency across topics
Within each model, we compared measures of consistency across topics. Fine-tuned models are much more inconsistent than base models when compared by topic. For example, llama3-base is about 60% more topic consistent than llama3. See Fig. LABEL:fig:models-topicwise-topic-consistency-hist-base. Namely, llama3 is significantly more inconsistent on euthanasia with a mean score of about .4 than it is on women’s rights with a mean of score of 0 while llama3-base is roughly as consistent in both cases (scoring about .2 and .1). See Fig. 1. In both topic and paraphrase consistency, fine-tuned models are more similar to our human participants in being inconsistently inconsistent (Fig. LABEL:fig:models-topicwise-topic-consistency-hist-base). For example, the mean topic inconsistency for our human respondents was .29 with a max of .44 and a min of 0, akin to the mean topic consistency of llama3 of .19 with a max of .45 and min of 0 compared to the mean for llama3-base of .12 with a max of .20 and min of .07.
@@ 6.2 Consistency by {un}controversial
Figure 7: Chat models are much less consistent on topics like euthanasia than they are for topics like women’s rights while base models are similarly consistent. Shown are the four topics with the highest (top row) and lowest (bottom row) topic inconsistency across models and human participants (n=84) in English on U.S.-based topics. Questions for each topic shown in Tab. 10 and 11.
@@
Comparing alignment fine-tuned models with their base model equivalents (Tab. 3), Fig. 6 shows that base models are more consistent compared to alignment fine-tuned models, especially on topic consistency. For example, llama3 is about 60% more topic inconsistent than llama3-base. While llama3 is about 33% less paraphrase consistent than llama3-base, all other chat models are more paraphrase inconsistent than their base models.
@@ 6.4 Consistency by use-case
We find that models are generally somewhat less consistent in the open-ended use-case than in the multiple-choice use-case (§3). This is more pronounced for yi and stability which are 27% and 57% more topic consistent on multiple-choice as shown in Fig. LABEL:fig:models-generation-consistency. Only llama2 is less topic consistent on multiple-choice with a reduction of 20%. Note that we use llama3 to judge the stance of the open-ended generations, and we find that it achieves substantial agreement with claude-3-opus and gpt-4o, with a median Fleiss’s Kappa of 0.7. (See Fig. 12.)
@@ 6.5 Can models be steered to certain values?
Figure 8: Chat models are somewhat less consistent in the open-ended use-case than in the multiple-choice use-case. We prompt gpt-4o, llama2, llama3 with U.S. topics and cmd-r, yi, and stability with German, Chinese, and Japanese topics, each in their respective dominant languages. We use llama3 to judge the stance of the open-ended generations. Error bars show 95% bootstrapped confidence intervals.fig:models-generation-consistency
@@
Figure 10: The five topics about which models and humans most disagreed for U.S.-based topics in English.
Scholars care about not just which values models express but also to which they are sensitive. Here we study whether models can be steered to answer in line with Schwartz’s values (Schwartz, 1992) as a proxy for value steerability in general. We choose Schwartz’s values because previous work is mixed on whether LLMs are steerable to them (Zhang et al., 2023; Yao et al., 2023; Fischer et al., 2023).
To determine whether prompting with certain value-words has any effect on models, we must first determine whether models can disambiguate between them. To do so, we prompted models with the questionnaire used to cluster and create Schwartz’s 11 values, the Portrait Values Questionnaire (PVQ-21). We then tested whether appending the name of each value (e.g. “universalism”) had a larger effect on the model response as compared to values unrelated to the question. (§A.4 offers a formal treatment. See §D.2 for an example.)
Which value was the most influential, the relevant value or an unrelated value? A rank of 0 indicates all of the unrelated values had a bigger effect than the related value while a rank of 11 (for the 12 values) means that the relevant value had a bigger effect than the unrelated values. While we would expect high rankings—high “steerability”—instead we find that unrelated values are more influential than relevant ones (Fig. 9). This means that the models were not steerable to these values.
@@ 7 Discussion
Prior work has argued that models either do (Durmus et al., 2024; Santurkar et al., 2023) or do not (Röttger et al., 2024; Shu et al., 2023) hold certain values. So: Are LLMs consistent over value-laden questions? While the answer is more yes than no, our findings show that the underlying complexity cannot be captured by a binary answer.
Indeed, unlike prior work (Röttger et al., 2024; Shu et al., 2023), we have found that large models ($>=34b$) are relatively consistent across our measures, performing on par with human participants on topic and paraphrase consistency (Fig. 4). Nonetheless, models’ consistency is not uniform.
In general, base models are more consistent than their fine-tuned counterparts (Fig. 6). Moreover, base models are more consistently consistent than fine-tuned ones. For example, llama3, like our human participants, is very consistent on women’s rights but very inconsistent on euthanasia while llama3-base does not exhibit such patterns (Fig. LABEL:fig:models-topicwise-topic-consistency-hist-base). Models are more consistent over uncontroversial than controversial questions (Fig. 5). We also measure how well models can be steered to particular values (§6.5), showing that models cannot be steered using a common set of values (Fig. 9).